OpenStudy (anonymous):

A genius society requires an IQ that is in the top 2% of the population in order to join. If an IQ test has a mean of 100 and a standard deviation of 15, what is the minimum qualifying score to join the genius society?

OpenStudy (anonymous):

we need Pr(X > x) = 0.02 which is like: pr(X<x) = 0.98 We will determine z: pr(Z<z) = 0.98 and in place of the table, I will just type it =norminv(0.98,0,1) which gives the number z satisfying Pr(Z<z) = 0.98 Result is 2.053748176 so x - 100 ▬▬▬..=...2.053748176 15 x = 2.053748176(15) + 100 = 130.8

OpenStudy (anonymous):

That is correct. The answer is 130. I just got my score back from test result and it is 130. :)